Use the following information for the month of July for questions 33-36 ABC Inc. uses the FIFO cost method in a perpetual inventory system. Jul 1 Beginning inventory 20 units @ $19 per unit Jul 7 Purchases Jul 8 Sales Jul 9 Sales Jul 10 Purchases Jul 22 Sales 70 units @ $20 per unit 50 units 25 units 50 units @ $22 per unit 40 units 33. The cost of goods sold for the July 8 sale was (a) $950. (b) $980 (c) $989. (d) $1,000

media%2Fc8e%2Fc8e9dd8b-e8f1-43a7-9039-6f
Answer #1
33
Cost of goods sold for Jul 8 = (20*19)+(30*20)= $980
34
Cost of goods sold for Jul 9 = (25*20)= $500
35
Cost of goods sold for July = 980+500+(15*20)+(25*22)= $2330
36
Ending inventory = 25*22= $550